# Open-Ended 3D Point Cloud Instance Segmentation

### v1

@@ Abstract
Open-Vocab 3D Instance Segmentation methods (OV-3DIS) have recently demonstrated their ability to generalize to unseen objects. However, these methods still depend on predefined class names during testing, restricting the autonomy of agents. To mitigate this constraint, we propose a novel problem termed Open-Ended 3D Instance Segmentation (OE-3DIS), which eliminates the necessity for predefined class names during testing. Moreover, we contribute a comprehensive set of strong baselines, derived from OV-3DIS approaches and leveraging 2D Multimodal Large Language Models. To assess the performance of our OE-3DIS system, we introduce a novel Open-Ended score, evaluating both the semantic and geometric quality of predicted masks and their associated class names, alongside the standard AP score. Our approach demonstrates significant performance improvements over the baselines on the ScanNet200 and ScanNet++ datasets. Remarkably, our method surpasses the performance of Open3DIS, the current state-of-the-art method in OV-3DIS, even in the absence of ground-truth object class names.
@@ 1 Introduction
3D point cloud instance segmentation (3DIS) chen2021hierarchical ; Schult23mask3d ; ngo2023isbnet ; he2021dyco3d ; vu2022softgroup ; jiang2020pointgroup , also known as closed-vocabulary 3D instance segmentation, aims to segment all points in a point cloud into instances of classes predefined in the training set. However, this approach is less practical for scenarios where the test classes are unknown or different from the training classes. This limitation has led to the development of open-vocabulary 3D instance segmentation (OV-3DIS) nguyen2023open3dis ; takmaz2023openmask3d ; lu2023ovir ; yin2023sai3d ; yan2024maskclustering ; Huang2023Segment3D . Despite these advancements, OV-3DIS methods face several practical challenges. One major challenge is that class names must be predefined during testing, necessitating human intervention for scene understanding. This requirement significantly impedes the perception of truly autonomous agents. A potential solution is to predefined large and comprehensive vocabularies; however, this can lead to inaccuracies and significantly degrade the performance of OV-3DIS when an excessive number of vocabularies are used.
To overcome these limitations, we introduce a novel task called Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS). Unlike traditional methods, OE-3DIS does not require predefined class names during testing. Given a 3D point cloud and RGBD sequence, the system automatically generates a set of 3D masks along with their class names. To evaluate the performance of OE-3DIS methods, we propose a new metric called the OE score, in addition to the standard AP score. The OE score measures both the intersection-over-union (IoU) of the predicted instance masks with the ground-truth masks and the semantic similarity between the predicted and ground-truth class names.
To address this new and challenging task, we investigate several typical approaches: leveraging OV-3DIS methods and utilizing Multimodal Large Language Models (MLLMs). In the first approach, a list of class names is either predefined from large vocabularies or pre-extracted from image taggers. In the second approach, 2D visual tokens extracted from a pretrained 2D CLIP encoder are lifted to 3D and used as 3D visual token inputs to an LLM to predict class names. Among these methods, we find that pointwise visual token lifting to 3D for use as input to the LLM performs the best and is selected as our chosen approach. This method performs on par with state-of-the-art OV-3DIS approaches, where ground-truth classes are provided. Additionally, this approach is training-free, leveraging a pretrained 2D vision encoder (e.g., CLIP radford2021learning ) and a pretrained LLM (e.g., Vicuna chiang2023vicuna ), thereby mitigating the drawback of limited training data faced by many 3D-LLM methods.
@@
To evaluate the performance of these methodologies, we conduct experiments on open-ended versions of two prominent 3D instance segmentation (3DIS) datasets: ScanNet200 rozenberszki2022language and ScanNet++ yeshwanthliu2023scannetpp . The results underscore the efficacy of our chosen approach over alternative baselines, demonstrating performance levels comparable to OV-3DIS methods that rely on ground-truth class names. Specifically, in ScanNet200, our approach attained an AP of 16.0, contrasting with the 22.2 AP achieved by Open3DIS, currently recognized as the state-of-the-art in OV-3DIS. However, our approach is superior in ScanNet++, where it outperforms Open3DIS by a significant margin (18.4 vs. 13.1).
@@
    We introduce the OE score, which evaluates both the IoU of predicted instance masks with ground-truth masks and the semantic similarity of predicted and ground-truth class names.
@@
    We explore approaches for OE-3DIS, including leveraging OV-3DIS methods and Multimodal Large Language Models (MLLMs).
  4. 4.
    We present a training-free OE-3DIS method that lifts 2D visual tokens to 3D for LLM input, utilizing pretrained models like EVA-CLIP and Vicuna. This approach outperforms other baselines and achieves performance comparable to OV-3DIS methods using ground-truth class names on ScanNet200 and ScanNet++ benchmarks.
@@ 2 Related Work
3D instance segmentation (3DIS) methods such as Mask3D Schult23mask3d , ISBNet ngo2023isbnet , PointGroup jiang2020pointgroup , and others lu2023query ; robert2023spt ; robert2024scalable ; Al_Khatib_2023_spatial cluster a point cloud scene into 3D instance masks of classes predefined in the training set. These methods utilize a 3D Convolutional backbone choy2019fully ; spconv2022 ; choy2020high to extract semantic information from the 3D scene. Subsequently, they employ either Dynamic Convolution-based he2021dyco3d or Grouping-based vu2022softgroup modules to generate 3D instance masks. Recently, some approaches have adopted techniques to back-project 2D information aggregated from multiple views onto the 3D point cloud to create an ensemble of 3D point cloud features puy24scalr ; Huang2023Segment3D ; Peng2023OpenScene ; guo2023samgraph . These 2D-derived features contain rich semantic information, while those derived from 3D capture the geometrical structure of 3D objects. Combined, they supervise a 3D instance decoder to refine segmentation masks. However, these methods are closed-vocab or cannot segment new classes in testing, limiting the capability of agents to understand new 3D scenes.
@@
Open-vocabulary 3D instance segmentation (OV-3DIS) aims at segmenting 3D objects of classes newly provided in testing. To provide 3D proposals for object recognition, OpenMask3D takmaz2023openmask3d and Lowis3D ding2023lowis3d employ 3DIS networks ngo2023isbnet ; Schult23mask3d to generate class-agnostic 3D proposals, while SAI3D yin2023sai3d , MaskClustering yan2024maskclustering , and OVIR lu2023ovir utilize 2D segmenter for producing masks for each view and consistently lift these masks to 3D. While 3DIS networks excel in capturing large geometrical structures, they often struggle to detect rare and small-shaped objects. Conversely, 2D segmenters are adept at focusing on small regions but face challenges in maintaining object consistency when lifting to the 3D point cloud. Open3DIS nguyen2023open3dis addresses these limitations by combining both 3D and 2D branches, resulting in superior results. This approach effectively captures rare and small objects while preserving the 3D geometrical structures of large objects using superpoint-level masks. While OV-3DIS is useful in some scenarios, the constraint of a predefined vocabulary set in testing requires human intervention, hindering truly autonomous agents.
@@
3D scene understanding with Large Language Models (LLMs). Utilizing LLMs for 3D scene understanding focuses on how objects are aligned, their directions, and their locations based on textual questions within 3D environments. This approach emphasizes the spatial aspects of language understanding in three dimensions of data. Previous works 3dllm ; huang2023embodied ; scanqa ; achlioptas2020referit_3d ; chen2020scanrefer have contributed to providing 3D spatial data with language for various applications, including 3D instance and scene captioning chen2023ll3da ; chen2023end ; chen2021scan2cap ; SpaCap3D ; huang2023embodied ; 3dvista ; 3dllm , 3D visual answering questions huang2023embodied ; scanqa ; ma2022sqa3d ; parelli2023clip ; 3dvista ; 3dllm ; delitzas2023multi ; chen2023ll3da , 3D visual grounding 3dvista ; 3dllm ; huang2023embodied ; achlioptas2020referit_3d ; chen2020scanrefer ; wu2022eda and supporting embodied AI tasks like planning and reasoning 3dllm ; 3dvista ; huang2023embodied ; chen2023ll3da .
@@
Open-ended 2D Image Understanding is an emerging task that addresses the need to recognize objects without predefined class names during training or testing. There is scant work on this task, with existing research primarily focusing on image classification nxtp ; conti2023vocabularyfree ; zhang2023recognize , object detection lin2024generateu , and instance segmentation you2023ferret ; yu2023towards ; Osprey . Standard 2D Multimodal LLMs (2D MLLMs), such as LLAVA liu2023llava , consist of a frozen vision encoder, a projector, and an LLM module. These models typically finetune either (1) the linear projector and the LLM or (2) a complex Q-Former projector. However, applying these methods for 3D scene understanding (3D-LLMs) is challenging due to the lack of sufficient 3D data and text description pairs to effectively train 3D-LLMs. In this paper, we present a novel approach to leveraging pretrained 2D MLLMs for OE-3DIS.
@@
OmniScient Model (OSM) yu2023towards is a recently proposed pretrained 2D Multimodal LLM for Open-ended 2D Instance Segmentation, which serves as the foundation for our proposed method. OSM comprises three main modules: a visual encoder, a MaskQ-Former, and a Large Language Model (LLM). The visual encoder is a pretrained EVA-CLIP fang2023eva , a variant of the CLIP model radford2021learning , which extracts high-resolution visual features using a sliding-window scheme and incorporates global positional embeddings to preserve spatial information. The MaskQ-Former, a customized version of Q-Former instructblip designed to focus on the mask region rather than the entire image, converts visual features into fixed-length visual tokens. These tokens are then input into the Vicuna LLM chiang2023vicuna . The LLM processes these tokens to answer the question, “What is in the segmentation mask?” by outputting the object name.
@@
In OSM, only the MaskQ-Former is trained to align visual features with the visual tokens for the LLM, while the visual encoder and LLM remain unchanged. The training datasets are large, including COCO coco , LVIS gupta2019lvis , ADE20K ade , and Cityscapes cordts2016cityscapes . This setup demonstrates a strong capability for recognizing objects without predefined class names in 2D images, inspiring us to extend this approach to 3D scene understanding.
@@ 3.1 Problem Statement
Given a 3D point cloud scene $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points with $xyz$ coordinates and associated $rgb$ colors, along with $T$ RGB-D frames with color images $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and depth ones $\{\mathbf{D}_{t}\}_{t=1}^{T}$, where $\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}_{t}\in\mathbb{R}_{+}^{H\times W}$, we aim to segment all $K$ object binary masks $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\mathbf{m}_{k}\in\{0,1\}^{N}$ and their associated class names $\{l_{k}\}_{k=1}^{K}$ without giving any predefined class names in testing.
The auxiliary information includes the intrinsic matrix $\mathbf{\Gamma}\in\mathbb{R}^{3\times 3}$ and the extrinsic matrix $\mathbf{[R|v]}_{t}\in\mathbb{R}^{3\times 4}$, where $H,W$ are the image’s height and width, $\mathbf{R}$ is a 3D rotation matrix, and $\mathbf{v}$ is a 3D translation vector. This composite matrix of rotation and translation converts coordinates from the global frame (of the point cloud) to the camera’s frame at view $t$.
@@ 3.2 Evaluation Metrics
To evaluate open-ended object detection or instance segmentation, where predicted class names may be similar but not exactly the same as ground-truth (GT) class names, prior work Rewatbowornwong2023ZeroGuideSeg proposed a label reassignment technique. This method uses text encoders (e.g., CLIP radford2021learning , BERT devlin2018bert , Sentence Transformer sentence-transformer to encode both the predicted and GT class names for each scene. It then matches each predicted class name to its closest GT class name based on cosine similarity. After this matching, the standard AP score is used to evaluate performance.
However, the AP score described above is based on a many-to-one matching where multiple predicted class names can be matched to the same ground-truth class name, leading to inaccuracies in fine-grained class prediction. To address this, we propose a new metric called the Open-ended score (OE score), which considers both textual semantic similarity and spatial similarity of predicted and ground-truth masks and their class names. Specifically, given $K$ predicted instances with masks, class names, and confidence scores $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\{l_{k}\}_{k=1}^{K}$, and $\{\rho_{k}\}_{k=1}^{K}$, and $J$ ground-truth instances with masks and class names as $\{\mathbf{m}_{j}\}_{j=1}^{J}$ and $\{l_{j}\}_{j=1}^{J}$, we first compute the cosine similarity $s_{k,j}$ between the text embeddings of the predicted and ground-truth class names, along with the IoU $o_{k,j}$ between predicted and ground-truth masks. Next, we construct the cost matrix $\mathbf{C}\in[0,1]^{K\times J}$ for one-to-one Hungarian matching as follows:
$\mathbf{C}(k,j)=-\sqrt[3]{o_{k,j}\cdot\rho_{k}\cdot\max(0,s_{k,j})}.$ (1)
In this way, only the best-matched 3D predictions in both text and mask are considered for evaluation. The final OE score is obtained by averaging all costs of matched pairs, or
OE $\text{OE score}=100*\frac{\sum_{(k,j)\in\text{all matched pairs}}-\mathbf{C}(k,j)}{J}.$ (2)
We average over $J$ GT masks since the number of GT masks is always smaller than the number of predicted masks.
@@ 3.3 Proposed Baselines
Since OE-3DIS is very new and challenging, even more so than OV-3DIS, we focus our efforts on investigating prominent baselines. These baselines are illustrated in Fig. 1. They require a list of class-agnostic 3D mask proposals pre-extracted from Open3DIS nguyen2023open3dis with the DETIC 2D segmenter.
Large-vocab approach (Fig. 1 - Left): We start with a simple OE-3DIS baseline by using a large vocabulary of 21K common classes introduced by DETIC detic as predefined class names for OV-3DIS methods like Open3DIS nguyen2023open3dis and OpenMask3D takmaz2023openmask3d . However, this approach fails to achieve robust class prediction. This is because the fixed large vocabulary set contains multiple synonyms, resulting in uninformative class predictions after the Softmax operation.
Image Tagging approach (Fig. 1 - Middle): To reduce the number of classes, we leverage image-tagging techniques such as RAM++ zhang2023recognize to obtain only relevant class names per scene. Specifically, for each input view, a set of image tags is generated and then combined across all processed input views. The resulting unified tag set serves as the vocabulary for OV-3DIS. However, these methods often produce inconsistent class names across views, leading to redundant and similar class names.
Maskwise approach (Fig. 1 - Right): To tackle the inconsistency in class names across views, we first apply an OE-2DIS method, such as OSM yu2023towards , to each view to obtain a list of 2D masks and their predicted fixed-length visual tokens. For each 3D mask proposal, we project it onto a view and associate it with the best-matched 2D mask based on IoU to obtain its 2D fixed-length visual tokens. The 3D visual tokens for the 3D mask are then aggregated by averaging the 2D visual tokens across views, which are subsequently input into a pretrained LLM to obtain the final class names. However, this approach heavily relies on the matching process between the 3D proposal and 2D masks in each view, which is often misaligned due to imperfect segmentation and depth maps. In other words, a 3D mask can project onto multiple 2D masks, and selecting only the best-matched 2D mask might result in the loss of important information.
Figure 2: Overview of our approach. First, we generate class-agnostic 2D instance segmentation masks for all views using segmenters like DETIC detic and SAM sam , and lift these 2D masks into 3D masks using Open3DIS nguyen2023open3dis . Simultaneously, the 2D masks and their corresponding RGB images are used to extract 2D visual tokens from an MLLM like OSM yu2023towards , which are then lifted into pointwise 3D visual tokens. Finally, for each 3D proposal mask, we aggregate the pointwise 3D visual tokens to form the final tokens for input to the LLM to predict final class names.

### v2

@@ Abstract
Open-vocabulary 3D Instance Segmentation methods (OV-3DIS) have recently demonstrated their generalization ability to unseen objects. However, these methods still depend on predefined class names during inference, restricting agents’ autonomy. To mitigate this constraint, we propose a novel problem termed Open-Ended 3D Instance Segmentation (OE-3DIS), which eliminates the necessity for predefined class names during testing. We present a comprehensive set of strong baselines inspired by OV-3DIS methodologies, utilizing 2D Multimodal Large Language Models. In addition, we introduce a novel token aggregation strategy that effectively fuses information from multiview images. To evaluate the performance of our OE-3DIS system, we benchmark both the proposed baselines and our method on two widely used indoor datasets: ScanNet200 and ScanNet++. Our approach achieves substantial performance gains over the baselines on both datasets. Notably, even without access to ground-truth object class names during inference, our method outperforms Open3DIS, the current state-of-the-art in OV-3DIS. Source code available at: https://github.com/PhucNDA/OE-3DIS.
Figure 1: During testing, Open-Vocabulary 3D Instance Segmentation (OV-3DIS) utilizes CLIP to generate confidence scores for provided text prompts. However, its reliance on a predefined vocabulary during inference necessitates human intervention, restricting autonomy in intelligent agents. In contrast, our proposed Open-Ended 3D Instance Segmentation (OE-3DIS) leverages Large Language Models to enable open interaction with humans, recognizing fine-grained objects without being constrained by predefined class labels.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
3D point cloud instance segmentation (3DIS) [5, 50, 37, 25, 52, 30], also known as closed-vocabulary 3D instance segmentation, aims to segment all points in a point cloud into instances of classes predefined in the training set. However, this approach is less practical for scenarios where the test classes are unknown or different from the training classes. This limitation has led to the development of open-vocabulary 3D instance segmentation (OV-3DIS) [39, 51, 35, 57, 55, 28]. Despite these advancements, OV-3DIS methods face several practical challenges. One major challenge is that class names must be predefined during inference, necessitating human intervention for scene understanding. This requirement significantly impedes the perception of truly autonomous agents. A potential solution is to predefined large and comprehensive vocabularies; however, this can lead to inaccuracies and significantly degrade the performance of OV-3DIS when an excessive number of vocabularies are used.
To overcome these limitations, we introduce a novel task called Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS) see Fig. 1. Unlike traditional methods, OE-3DIS does not require predefined class names during testing. Given a 3D point cloud and RGBD sequence, the system automatically generates a set of 3D masks along with their class names. To evaluate the performance of OE-3DIS methods, we follow [46] to leverage an adaptive protocol of standard AP score calculation. To tackle this new and challenging task, we introduce a novel method along with our proposed three established baselines that leverage OV-3DIS techniques and Multimodal Large Language Models (MLLMs). In the first two baselines, class names are either predefined from extensive vocabularies or extracted using image taggers. The third baseline employs 2D visual tokens, obtained from a pretrained 2D CLIP encoder, which are then processed through a Visual Token Aggregation mechanism before being fed into an LLM to predict class names. Finally, our proposed method takes this further by consistently aggregating 2D visual tokens from multiview images, transforming them into dense 3D point cloud tokens that enable real-time querying. Notably, our pointwise visual token lifting approach, which feeds these tokens into an LLM, achieves the best performance across multiple benchmarks, matching state-of-the-art OV-3DIS methods that rely on ground-truth class labels. Furthermore, both our proposed method and designed baselines are entirely training-free, utilizing only pretrained 2D vision encoders (e.g., CLIP [43]) and pretrained LLMs (e.g., Vicuna [10]). The training-free approach effectively mitigates the prevalent issue of insufficient training data that hampers many existing 3D-LLM techniques.
@@
To evaluate the performance of these methods, we conduct experiments on open-ended versions of two prominent 3D instance segmentation (3DIS) datasets: ScanNet200 [49] and ScanNet++ [56]. The results underscore the efficacy of our chosen approach over alternative baselines, demonstrating performance levels comparable to OV-3DIS methods that rely on ground-truth class names. Specifically, in ScanNet200, our approach attained an AP of 16.0, contrasting with the 22.2 AP achieved by Open3DIS, currently recognized as the state-of-the-art in OV-3DIS. However, our approach is superior in ScanNet++, where it outperforms Open3DIS by a significant margin (18.4 vs. 13.1 in AP).
@@
    We establish solid baselines for OE-3DIS, including leveraging OV-3DIS methods and Multimodal Large Language Models (MLLMs).
@@
    We present a training-free OE-3DIS method that lifts 2D visual tokens to 3D and utilizes pretrained Multimodal LLMs to output the final object classes.
@@ 2 Related Work
3D instance segmentation (3DIS) methods such as Mask3D [50], ISBNet [37], PointGroup [30], and others [34, 47, 48, 2] cluster a point cloud scene into 3D instance masks of classes predefined in the training set. These methods utilize a 3D Convolutional backbone [11, 15, 12] to extract semantic information from the 3D scene. Subsequently, they employ either Dynamic Convolution-based [25] or Grouping-based [52] modules to generate 3D instance masks. Recently, some approaches have adopted techniques to back-project 2D information aggregated from multiple views onto the 3D point cloud to create an ensemble of 3D point cloud features [42, 28, 41, 23]. These 2D-derived features contain rich semantic information, while those derived from 3D capture the geometrical structure of 3D objects. Combined, they supervise a 3D instance decoder to refine segmentation masks. However, these methods are closed-vocab or cannot segment new classes in testing, limiting the capability to understand new 3D scenes.
@@
Open-vocabulary 3D instance segmentation (OV-3DIS) aims at segmenting 3D objects of classes newly provided in testing. To provide 3D proposals for object recognition, OpenMask3D [51] and Lowis3D [21] employ 3DIS networks [37, 50] to generate class-agnostic 3D proposals, while SAI3D [57], MaskClustering [55], OVIR [35] and Any3DIS [38] utilize 2D segmenter for producing masks for each view and lift these masks to 3D. While 3DIS networks excel in capturing large geometrical structures, they often struggle to detect rare and small-shaped objects. Conversely, 2D segmenters are adept at focusing on small regions but face challenges in maintaining object consistency when lifting to the 3D point cloud. Open3DIS [39] addresses these limitations by combining both 3D and 2D branches, resulting in superior results. This approach effectively captures rare and small objects while preserving the 3D geometrical structures of large objects using superpoint-level masks. While OV-3DIS is useful in some scenarios, the constraint of a predefined vocabulary set in inference requires human intervention, hindering very autonomous agents.
@@
3D scene understanding with Large Language Models (LLMs). Utilizing LLMs for 3D scene understanding focuses on how objects are aligned, their directions, and their locations based on textual questions within 3D environments. This approach emphasizes the spatial aspects of language understanding in three dimensions of data. Previous works [26, 27, 3, 1, 4] have contributed to providing 3D spatial data with language for various applications, including 3D instance and scene captioning [6, 7, 8, 53, 27, 64, 26], 3D visual answering questions [27, 3, 36, 40, 64, 26, 19, 6], 3D visual grounding [64, 26, 27, 1, 4, 54] and supporting embodied AI tasks like planning and reasoning [26, 64, 27, 6].
@@
Open-ended 2D Image Understanding is an emerging task that addresses the need to recognize objects without predefined class names during training or testing. There is scant work on this task, with existing research primarily focusing on image classification [61, 14, 29], object detection [13], and instance segmentation [58, 59, 60]. Standard 2D Multimodal LLMs (2D MLLMs), such as LLAVA [33], consist of a frozen vision encoder, a projector, and an LLM module. These models typically finetune either (1) the linear projector and the LLM or (2) a complex Q-Former projector. However, applying these methods for 3D scene understanding (3D-LLMs) is challenging due to the lack of sufficient 3D data and text description pairs to effectively train 3D-LLMs. In this paper, we present a novel approach to leveraging pretrained 2D MLLMs for OE-3DIS.
@@
OmniScient Model (OSM) [59] is a recently proposed pretrained 2D Multimodal LLM for Open-ended 2D Instance Segmentation, which serves as the foundation for our proposed method. OSM comprises three main modules: a visual encoder, a MaskQ-Former, and a Large Language Model (LLM). The visual encoder is a pretrained EVA-CLIP [22], a variant of the CLIP model [43], which extracts high-resolution visual features using a sliding-window scheme and incorporates global positional embeddings to preserve spatial information. The MaskQ-Former, a customized version of Q-Former [18] designed to focus on the mask region rather than the entire image, converts visual features into fixed-length visual tokens. These tokens are then input into the Vicuna LLM [10]. The LLM processes these tokens to answer the question, “What is in the segmentation mask?” by outputting the object name.
@@
In OSM, only the MaskQ-Former is trained to align visual features with the visual tokens for the LLM, while the visual encoder and LLM remain unchanged. The training datasets are large, including COCO [32], LVIS [24], ADE20K [62], and Cityscapes [16]. This setup demonstrates a strong capability for recognizing objects without predefined class names in 2D images, inspiring us to extend this approach to 3D scene understanding.
@@ 3.1 Problem Statement
Given a 3D point cloud scene $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points with xyzxyz coordinates and associated rgbrgb colors, along with $T$ RGB-D frames with color images $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and depth ones $\{\mathbf{D}_{t}\}_{t=1}^{T}$, where $\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}_{t}\in\mathbb{R}_{+}^{H\times W}$, we aim to segment all $K$ object binary masks $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\mathbf{m}_{k}\in\{0,1\}^{N}$ and their associated class names $\{l_{k}\}_{k=1}^{K}$ without giving any predefined class names in inference. Of course, we do need GT class names during evaluation to assess our method’s performance.
The information includes the intrinsic $\mathbf{\Gamma}\in\mathbb{R}^{3\times 3}$ and the extrinsic $\mathbf{[R|v]}_{t}\in\mathbb{R}^{3\times 4}$. In this context, $H$ and $W$ represent the height and width of the image, respectively. The matrix $\mathbf{R}$ is a 3D rotation matrix, while $\mathbf{v}$ is a 3D translation vector. This composite matrix, which combines rotation and translation, converts coordinates from the global frame of the point cloud to the camera’s frame at view $t$.
@@ 3.2 Evaluation Metrics
To evaluate open-ended object detection or instance segmentation, where predicted class names may be similar but not exactly the same as ground-truth (GT) class names, prior work [46] proposed a label reassignment technique. This method uses text encoders (e.g., CLIP [43], BERT [20], Sentence Transformer [44] to encode both the predicted and GT class names for each scene. It then matches each predicted class name to its closest GT class name based on cosine similarity. After this matching, the standard AP score is used to evaluate performance.
@@ 3.3 Proposed Baselines
Since OE-3DIS is very new and challenging, we focus our efforts on investigating prominent baselines. These baselines are illustrated in Fig. 2. They require a list of class-agnostic 3D mask proposals pre-extracted from Open3DIS [39] with the DETIC 2D segmenter.
Large-vocab approach (Fig. 2 - Left): We start with a simple OE-3DIS baseline by using a large vocabulary of 21K common classes introduced by DETIC [63] as predefined class names for OV-3DIS methods like Open3DIS [39] and OpenMask3D [51]. However, this approach fails to achieve robust class prediction. This is because the fixed large vocabulary set contains multiple synonyms, resulting in uninformative class predictions after the Softmax operation.
Image Tagging approach (Fig. 2 - Middle): To reduce the number of classes, we leverage image-tagging techniques such as RAM++ [29] to obtain only relevant class names per scene. Specifically, for each input view, a set of image tags is generated and then combined across all processed input views. The resulting unified tag set serves as the vocabulary for OV-3DIS. However, these methods often produce inconsistent class names across views, leading to redundant and similar class names.
Maskwise approach (Fig. 2 - Right): To tackle the inconsistency in class names across views, we first apply an OE-2DIS method, such as OSM [59], to each view to obtain a list of 2D masks and their predicted fixed-length visual tokens. For each 3D mask proposal, we project it onto a view and associate it with the best-matched 2D mask based on IoU to obtain its 2D fixed-length visual tokens. The 3D visual tokens for the 3D mask are then aggregated by averaging the 2D visual tokens across views, which are subsequently input into a pretrained LLM to obtain the final class names. However, this approach relies on matching 3D proposals with 2D masks, which is often misaligned due to segmentation and depth map imperfections. A 3D mask can project onto multiple 2D masks, and selecting only the best match may discard valuable information.
Figure 3: Overview of our approach. First, we generate class-agnostic 2D instance segmentation masks for all views using segmenters like DETIC [63] and SAM [31], and lift these 2D masks into 3D masks using Open3DIS [39]. Simultaneously, the 2D masks and their corresponding RGB images are used to extract 2D visual tokens from an MLLM like OSM [59], which are then lifted into pointwise 3D visual tokens. Finally, for each 3D proposal mask, we aggregate the pointwise 3D visual tokens to form the final tokens for input to the LLM to predict final class names.